# Post-detection inference for sequential changepoint localization

### v2

@@ 2.4 Unconditional coverage when $\mathcal{A}$ controls PFA
When $T\ll$ ARL, we typically have $\mathbb{\mathbb{P}}_{\infty}(\tau<T)\approx 0$ and so $\mathbb{P}_{T}(T\in\mathcal{C})\gtrsim 1-\alpha$. The confidence set constructed using our method will approximately achieve the desired unconditional coverage level $1-\alpha,$ in situations where the probability of false detections before the true changepoint is negligible. However, $T$ is not known to us, and hence, we are unable to produce a suitable bound on the unconditional coverage for the detection algorithms that almost surely stop at a finite time, even when there is no change.
2.5 Unconditional coverage when $\mathcal{A}$ has finite ARL
If the detection algorithm has finite ARL (i.e., $\tau$ is finite almost surely even if there is no change), then it is impossible for any confidence set $\mathcal{C}^{*}$, which is a subset of $\{1,\cdots,\tau\}$, to produce an unconditional coverage guarantee at a predetermined level $1-\alpha$. The intuitive reason is that if $\tau$ is finite almost surely, it is possibly a false detection, i.e., $\tau<T$, in which case the set $\{1,\cdots,\tau\}$ itself does not cover $T$. The next proposition formalizes this fact.
Proposition 2.4.
If a confidence set $\mathcal{C}^{*}$ that is a subset of $\{1,\cdots,\tau\}$ satisfies an unconditional coverage guarantee $\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha$, for some $\alpha\in(0,1)$, then $\mathbb{P}_{\infty}(\tau=\infty)\geq 1-\alpha$ (meaning that the PFA of the algorithm is at most $\alpha$ and thus the ARL of the algorithm is infinite). Said differently, for any algorithm with finite ARL, it is impossible to construct a valid confidence set that is only a subset of $\{1,\dots,\tau\}$.
The proposition says that if we have a confidence set with high unconditional coverage, then the detection algorithm never stops with high probability if there is no change.
Therefore, in order to achieve an unconditional coverage guarantee with $\mathcal{A}$ having finite ARL, we must extend our confidence set beyond the stopping time $\tau$. So, we consider the union of the confidence set constructed using Algorithm 1 and all integers greater than $\tau$. The following proposition establishes that this augmented confidence set achieves the desired coverage guarantee of $1-\alpha$.
Proposition 2.5.
Let $\mathcal{C}$ be the confidence set for changepoint $T\in\mathbb{N}$ produced by Algorithm 1, for some $\alpha\in(0,1)$, $B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$ and change detection algorithm $\mathcal{A}$. Then, $\mathcal{C}^{*}:=\mathcal{C}\cup\{\tau+1,\tau+2,\cdots\}$ satisfies
$\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha.$ (2.10)
So, the confidence set constructed using Algorithm 1 can be interpreted as the pre-detection part of the augmented confidence set, which has the desired coverage guarantee of $1-\alpha$.
2.6 Conditional coverage when $\mathcal{A}$ has finite ARL
In the previous subsection, we established that the conventional notion of coverage is unsuitable for detection algorithms with finite ARL, where the confidence set is a subset of the random set $\{1,\cdots,\tau\}$. Intuitively, it is more appropriate to consider the conditional coverage, given that no false detection occurs, i.e., $T\leq\tau$. This adjustment is crucial because discussing coverage becomes meaningless when $\tau<T$, meaning that the detection occurs before the true changepoint — the target point we wish to cover by our confidence set.
The following is an immediate corollary of Theorem 2.2, where we show a bound on the conditional coverage of the confidence set produced by Algorithm 1.
Corollary 2.6.
For any $\alpha\in(0,1)$, $B\in\mathbb{N}$ and $T\in\mathbb{N}$, and change detection algorithm $\mathcal{A}$ satisfying $\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}$ produced by Algorithm 1 satisfies
$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\alpha}{\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)}.$ (2.11)
Therefore, if $\mathcal{A}$ controls PFA at level $\delta\in(0,1)$,
$\mathbb{P}_{T}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{\alpha}{1-\delta}.$ (2.12)
The restriction that $\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)\neq 0$ is very weak and clearly necessary for conditional coverage. Since $T$ is arbitrary, the condition means that there is no $S<\infty$ satisfying $\mathbb{P}_{\infty}(\tau<S)=1$. If the algorithm always stops under $\mathbb{P}_{\infty}$ before $S$, then it can never detect any change that occurs after $S$. This is because the information that distinguishes $\mathbb{P}_{\infty}$ from $\mathbb{P}_{T}$ for $T>S$ occurs after $S$, but the algorithm has halted before $S$.
When $\tau$ is finite almost surely (i.e., PFA is $1$), we cannot have a positive lower bound on $\mathbb{\mathbb{P}}_{\infty}(\tau\geq T)$ uniformly over $T\in\mathbb{N}$. Therefore, Algorithm 1 does not guarantee conditional coverage at some prespecified level for detection algorithms with finite ARL (which implies that PFA is 1, although the converse is not true). Furthermore, if $\mathcal{A}$ is completely black-box — meaning its PFA is unknown — it becomes infeasible to construct a confidence set with the desired conditional or unconditional coverage using this approach. We propose a simple modification to address the issue in the following subsection, which applies to any black-box detection algorithm.
2.7 Modified confidence set when $\mathcal{A}$ has finite ARL (or $\mathcal{A}$ is a blackbox)
Equation 2.11 indicates that to ensure a conditional coverage guarantee of $1-\alpha$, we observe that we need to test $H_{0,T}$ at level $\alpha\times\mathbb{\mathbb{P}}_{\infty}(\tau\geq T),$ rather than simply $\alpha$. To achieve that, for each $t\in\{1,\cdots,\tau\}$, we test $H_{0,t}$ at level $\alpha\times\mathbb{\mathbb{P}}_{\infty}(\tau\geq t).$ For each $i\in\{1,\cdots,N\}$, we generate a sequence $\{Z_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0}$ until a change is detected, recording the detection times as $\tau_{i}$, which help to estimate the probability $\mathbb{P}_{\infty}(\tau\geq t)$, for $t=1,\cdots,\tau$. Therefore, it suffices to draw data points only up to $\min\{\tau,\tau_{j}\}$, for $j=1,\cdots,N.$
Then, we draw $B$ many independent streams of data under $H_{0,t}$. For the $j$th sequence, denoted as $\{X_{n}^{j}\}_{n}$, we simulate data points until a change is detected at the stopping time $\tau_{j}^{\prime}$, using the same change detection method $\mathcal{A}$. Then, we compute the test statistic $M_{t,L}^{j}$ as described in Section 2.2, based on $\tau_{j}^{\prime}\wedge L$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$. Our confidence set is the collection of all $t\in\{1,\cdots,\tau\}$, for which we fail to reject $H_{0,t}$, i.e.,
$\mathcal{C}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})\right\},$ (2.13)
where $\hat{\alpha}_{t}=\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$. Algorithm 2 contains a structured overview of the method.
Input: $\alpha\in(0,1)$, $N\in\mathbb{N},B\in\mathbb{N}$, change detection algorithm $\mathcal{A}$ and data $\{X_{n}\}_{n}$ until a change is detected at $\tau$ using $\mathcal{A}$
Output: A confidence set $\mathcal{C}$ for the changepoint $T$
1 $\mathcal{C}=\varnothing$
2 for $i=1,\cdots,N$ do
3       Simulate $\{Z_{n}^{i}\}_{n}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0}$, until $\mathcal{A}$ detects a change at $\tau_{j}$, or until $\tau$, whichever happens first.
4
5 end for
6for $t=1,\cdots,\tau$ do
7       Compute some test statistic $M_{t}$ (e.g. (2.9)) based on $\tau$ and $X_{1},\cdots,X_{\tau}$
8       for $j=1,\cdots,B$ do
9             Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+2}^{j}\cdots\stackrel{{\scriptstyle iid}}{{\sim}}F_{1}$, until $\mathcal{A}$ detects a change at $\tau_{j}^{\prime}$, or time $L$, whichever happens first.
10             Subroutine $\mathcal{L}$: Compute $M_{t,L}^{j}$ (defined in (2.3),(2.4)) based on $\tau_{j}^{\prime}$, $L$ and $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$
11
12       end for
13      Compute $\hat{\alpha}_{t}:=\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N$
14       if $M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})$ then
15            $\mathcal{C}=\mathcal{C}\cup\{t\}$
16       end if
17
18 end for
Algorithm 2 CI for known pre-change $F_{0}$ and post-change $F_{1}$
By suitably adjusting the test levels for each $t$, this approach ensures that the desired conditional coverage guarantee is achieved. Notably, this guarantee does not depend on any specific property of $\mathcal{A}$ (e.g., its PFA or ARL). We prove it in the following theorem.

### v3

@@ 4.3 Adaptive confidence set for parametric pre- and post-change classes
As a special case, when pre- and post-change parameters are known, i.e., $\Theta_{i}=\{\theta_{i}\}$ are singleton sets, $i=0,1$, we have $\mathcal{S}=\{\theta_{0}\},\mathcal{S}^{\prime}=\{\theta_{1}\}$ with $\beta=0$. So, in this case, Algorithms 3 and 2 reduces to our algorithm designed for the known pre- and post-change setting, and similarly, Theorems 4.4 and 4.5 reduces to Theorem 4.1.
5 Confidence set for pre- and post-change parameters
Beyond localizing the changepoint, another natural inferential goal is to estimate the pre- and post-change parameters, assuming that the pre- and post-change distributions follow some parametric composite models: $\mathcal{P}_{i}=\{F_{\theta}:\theta\in\Theta_{i}\};i=0,1$. While our primary focus is on constructing a confidence set $\mathcal{C}$ for the changepoint $T$, we now outline how one can derive confidence sets for $\theta_{0}$ and $\theta_{1}$ using $\mathcal{C}$.
The key idea is as follows: once a valid confidence set $\mathcal{C}\subseteq\{1,\cdots,\tau\}$ for $T$ is constructed, this implicitly identifies a subset of data that is likely to have been generated under the pre- and post-change regimes.
For estimating the post-change parameter, one natural idea might be to consider the samples starting from $\max\mathcal{C}$ till the detection time $\tau$. But since both starting and ending times are random here, this strategy does not yield valid coverage guarantees. Instead, we fix each $t\in\mathcal{C}$ and construct a confidence sequence based on $X_{t},\cdots,X_{\tau}$ for $\theta_{1}$, denoted as $\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*})$ (recall the definition in (4.3)), for some $\eta_{1}\in(0,1)$, whose value is less than the target non-coverage, $\alpha^{\prime}$. Finally, the union of these CS across $t\in\mathcal{C}$, yields the post-detection confidence interval for $\theta_{1}$:
$\cup_{t\in\mathcal{C}}\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*}).$ (5.1)
There is a natural tension between how quickly $\mathcal{A}$ detects a change from $\theta_{0}$ and the tightness of inference on $\theta_{1}$ — if $\mathcal{A}$ stops very quickly, we will have very few data points with which to estimate $\theta_{1}$, leading to wider intervals. This is not a drawback of our methodology, but an inherent limitation of any post-detection inference method.
Analogously, for the pre-change parameter $\theta_{0}$, we fix each $t\in\mathcal{C}$ and construct a confidence interval for $\theta_{0}$ based on the data ${X_{1},\ldots,X_{t-1}}$, denoted as $\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*})$, for some $\eta_{0}\in(0,1)$, chosen to be smaller than the target (conditional) non-coverage level, $\alpha^{\prime\prime}$. The final confidence set for $\theta_{0}$ is the following:
$\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*}).$ (5.2)
We formalize its coverage guarantee in the next theorem, which is nonasymptotic and agnostic to the detection algorithm, unlike existing methods [39, 36].